# Styx: Transactional Stateful Functions on Streaming Dataflows

### v2

@@ 4. Styx’s Architecture
In this section, we describe the components of Styx and the main design decisions we had to consider along the way as they appear in Figure 3.
@@ 4.1. Components
Coordinator. The coordinator manages and monitors the Styx workers. Its primary task is to manage the runtime state of the cluster (parallelism, dataflow state, partition locations, etc.) as well as perform scheduling and health monitoring. Currently, Styx implements round-robin scheduling, but we plan to add more complex schedulers in the future to be able to deal with load balancing. Styx monitors the cluster’s health using a typical heartbeat mechanism and initiates the fault-tolerance mechanism (Section 7) once a worker fails. Furthermore, it stores worker metrics that can be used by a dashboard to provide insights on the cluster, such as latency, the size of the sequences being processed, and partitioning imbalances.
@@
Worker. As depicted in Figure 3, the worker is the primary component of Styx, managing networking, processing, and application state. At its core, it consists of two primary coroutines. The first coroutine ingests messages for its assigned partitions in a push-pull fashion from durable queue systems, such as Kafka, and sequences them. The second coroutine receives a sequence of transactions and initiates the transaction processing. By utilizing the coroutine execution model, Styx increases its efficiency since the most significant latency factor is waiting for network or state-access calls. Coroutines allow for single-threaded concurrent execution, switching between coroutines when one gets suspended during a network call, allowing others to make processing progress. Once the network call is completed, the suspended coroutine can resume processing.
@@
Partitioning Stateful Entities Across Workers. Styx makes use of the entities’ key to distribute those entities and their associated state across a number of workers. By default, each worker is assigned a set of keys using hash partitioning.
@@
Replayable Source. An integral part of the architecture to support the fault tolerance scheme that Styx employs is a replayable source. Other than an ingress to the system providing the function calls from external systems (e.g., from a REST gateway API), the source must be able to deterministically replay messages based on an offset when a failure occurs. This is necessary to produce the same sequence of transactions after the recovery is complete and to enable early replies (Section 7). Currently, Styx uses Apache Kafka for that purpose.
@@
Durable Snapshot Store. Alongside the replayable source, durable storage is necessary for storing the workers’ snapshots. Currently, Styx uses Minio, an open-source S3 clone, to store the incremental snapshots as binary data files.
4.2. Function/Transaction Execution Pipeline
@@
In Styx, the terms function and transaction can be used interchangeably. A function triggers a call graph of function invocations within the system. Each such function can have its own state; the side-effects of the call graph of function invocations are transactionally committed to the state. In Figure 3, once a make_reservation function enters the system, it is persisted and replicated in the replayable source. Then, based on the partitioning, a worker ingests the function call into its sequencer and transforms it into a transaction by assigning a TID. The next step is to execute the transaction based on the transactional protocol detailed in Section 5 and Section 6. In the make_reservation case, the transaction consists of two functions: reserve_hotel and reserve_flight. For this example, let us assume that reserve_hotel is a local function call and reserve_flight a remote. reserve_hotel will execute locally (in an asynchronous fashion using co-routines) and apply state changes. In contrast, reserve_flight will execute asynchronously on a remote worker, applying state changes on the remote with zero state transfer other than the function parameters.
@@ 5. Sequencing & Function Execution
The deterministic execution of functions with serializable guarantees requires a sequencing step that assigns a sequence ID (TID), which, in combination with the read/write (RW) sets, can be used for conflict resolution (Section 6). The challenge that we tackle in this section is determining the boundaries of transactions (i.e., when transactional function executions begin and complete), which emerges from the execution of arbitrary function call graphs Section 5.3.
5.1. Function Sequencing
@@
In this section, we discuss the sequencing mechanism ( ) of Styx. Deterministic databases ensure the serializable execution of transactions by creating a global sequence. In Calvin (Thomson et al., 2012), the authors propose a partitioned sequencer that retrieves the global sequence by communicating across all partitions, performing a deterministic round-robin.
@@
Eliminating Sequencer Synchronization. Instead of the original sequencer of Calvin that sends messages for the deterministic round-robin, Styx adopts a method similar to Mencius (Mao et al., 2008), allowing Styx to acquire a global sequence without any communication between the sequencers (). This is achieved by having each sequencer assign unique transaction identifiers (TIDs) as follows:
@@
where is the sequencer id assigned by the Styx coordinator in the registration phase, is a local counter of each sequencer specifying how many TIDs it has assigned thus far and is the total number of sequencers in the Styx cluster. As an example in Figure 5 there are three sequencers with leading to the following sequences: , , . Thus, this arrangement leads to the same result as the deterministic round-robin schedule without any messaging between the sequencers.
In the example of Figure 4, the sequencers of the three workers will sequence , and to , and respectively.
Sequence Imbalance. A consideration with the approach used in Styx is the possibility of imbalance between the produced sequences. In the worst-case scenario where the total traffic in an epoch is directed to a single sequencer, the local counter will increment by that amount. Then, in the next epoch, that sequencer will produce larger TIDs than the other sequencers, and consequently, the new transactions it receives will have a lower priority for execution. In concurrency-related conflicts, this will increase latencies for that worker node. To avoid this issue, we pass all within the commit message so that each worker can set its local counter () as the of all, leading to balanced sequences in every epoch.
@@
Replication and Logging. There is no need for replication and logging of the sequence within Styx since the input is logged and replicated within the replayable source. After replaying in case of failure, the sequencers will produce the exact same sequence.
5.2. Call-Graph Discovery: Dry Run
@@
After sequencing, Styx needs to execute the sequenced transactions and determine their call graphs and read/write sets ( ). To this end, the function execution runtime ingests a given sequence of transactions to process in a given epoch. The number of transactions per epoch is either set by a timeout (by default, Styx ingests all transactions that have arrived every millisecond) or by a configurable maximum number of transactions that can run per epoch (by default 1000 transactions per epoch). We have chosen an epoch-based approach since processing the incoming transactions in a batch will lead to higher throughput.
@@
Styx’s runtime executes all the sequenced transactions on a snapshot of the data to retrieve the read/write sets. Transactions that span multiple workers will implicitly change the read/write sets of the remote workers via function calls. There is an additional issue related to discovering the read/write set of a transaction: before the functions execute, the call graph of the transaction and the remote function calls are unknown. This is an issue because the protocol requires all transactions to be completed before proceeding to the next phase. To tackle this problem, Styx proposes a function chaining protocol that is explained in more detail in Section 5.3.
@@
After this phase, all the stateful functions that comprise transactions would have finished execution and the read/write sets will be known. In Figure 4, transactions , , and will execute and create the following RW sets: , and and .
5.3. Function Acknowledgement Shares
@@
In the SFaaS paradigm, the call graph that forms a transaction is unknown; functions could be coded by different people, calling other functions written by others, and so on. This makes it difficult to determine when a transaction has finished processing and is necessary because phase can only start after all transactions have finished processing. To that end, each asynchronous function call of a given transaction is assigned an ack_share. A given function knows how many shares to create by counting the number of asynchronous function calls during its runtime. The caller function then sends the respective acknowledgment shares to the downstream functions. For instance, in Figure 6, the transaction entry-point (root of the tree) calls three remote functions, splitting the ack_share into three parts (3 x ). The left-most function invokes only one other function and passes to it its complete ack_share (). The middle function does not call any functions, so it returns the share to the transaction entry point (root), and the right-most function calls two other functions, splitting its share () to 2 x . After all the function calls are complete, the entry-point should have collected all the shares. When the sum of the received shares adds to 1, the root/entry-point function can safely deduce that execution is complete.
@@
This design was devised for two reasons: i) if every participating function just sent an ack when it was done, the root would not know how many acks to expect in order to decide whether the entire execution is finished, and ii) if we used floats instead of fractions we could stumble upon a challenge related to adding floating point numbers. For instance, if we consider floating point numbers in the aforementioned example of the three function calls, the sum of all shares would not equal 1, but 0.99 since each share contributes 0.33. Hence, we cannot accurately round inexact division numbers, and therefore, Styx makes use of fraction mathematics instead.
@@
A solution close to the ack_share is the one of distributed futures (Wang et al., 2021). However, it would not work in the SFaaS context because it either requires information about the entire call graph for it to work asynchronously, or it would need to create a chain of futures that would make it synchronous and hence it would introduce high latency for our use case.
@@ 6. Committing Transactions
After the completion of an epoch’s dry run, Styx needs to determine which transactions will commit and which will abort based on the transactions’ RW sets and TIDs. To this end, this section presents the two commit phases: an optimistic lock-free phase and a lock-based phase that leverages the RW sets to execute a commit phase by acquiring locks ordered by TID. To make the second phase faster, we have devised a call-graph caching scheme that can reuse the dry run’s call graph to avoid re-executing long call-graph chains (Section 6.3).
@@ 6.1. Lock-free Commit Phase
In case of conflict (i.e., a transaction writes a key that another transaction also reads or writes on) similarly to (Lu et al., 2020), only the transaction with the lowest transaction ID will succeed to commit ( ). The transactions that have not been committed are put in a queue to be executed in the next phase (maintaining their previously assigned ID). In addition, workers broadcast their local conflicts to every other worker: this way, every worker retains a global view of all the aborted/rescheduled transactions and can decide, locally, which transactions can be committed. Finally, note that transactions can also abort, not because of conflicts, but due to application logic causes (e.g., by throwing an exception). In that case, Styx removes the related entries from the read/write sets to reduce possible conflicts further.
In this phase, all the transactions that have not been part of a conflict apply their writes to the state and reply back to the clients. In the example shown in Figure 4, only can commit in and due to conflicts in the RW sets of regarding and ; more specifically in keys and .
@@ 6.2. Lock-based Commit Phase
In the previous phase, , only transactions without conflicts can be committed. We now explain how Styx deals with transactions that have not been committed in a given epoch due to conflicts ( ). First, Styx trivially acquires locks in a given sequence ordered by transaction ID. Then, it reruns all transactions concurrently since all the read/write sets are known and commits them. However, if a transaction’s read/write set changes in this phase, Styx aborts the transaction and recomputes its read/write set in the next epoch. Now, in Figure 4, can sequentially acquire locks for and , leading to their commits in and . In Section 6.3, we present an optimization for this phase that we introduced in Styx, which reduces its latency impact even further.
@@ 6.3. Call-Graph Caching
As depicted in Figure 4, the lock-based commit phase is used to execute any transactions that did not commit during the lock-free commit phase . Thus, by the time the lock-based commit phase is ready to execute, the state of the database may have changed since the lock-free commit. As a result, function invocations need to be re-executed to take into account the data updates.
On the left of Figure 7, we depict such a function invocation. At time , F is invoked, which in turn invokes two function chains: and . Once the two function chains finish their execution (on time and respectively), they can acknowledge their termination to the root call .
Potential for Caching. During our early experiments, we noticed cases where is invoked and the parameters with which it calls (and in turn the invocations across the call chain) do not change. The same applied to the RW set of those function invocations; the RW sets remained unchanged. Since Styx tracks those call parameters as well as the functions’ RW sets, Styx can simply cache input parameters during the lock-free commit phase and re-use them during the lock-based commit in order to avoid long sequential re-executions along the call chains. This case is depicted on the right of Figure 7: the function-call chain does not need to be invoked in a sequential manner from all the way to , leading to high-latency. Instead, the individual workers can simply re-invoke those function calls locally and in parallel. As a result, all functions can execute in parallel and save on latency and network overhead ( in Figure 7).
Conditions for Parallel Function Re-invocation. Intuitively, if the parameters with which e.g., is called, and the RW set of remains the same, we can safely assume that function can be invoked in parallel without having to be invoked sequentially by . If those functions are successfully completed and acknowledge their completion to the root function , it means that the transaction can be committed. To the contrary, if the RW set of any of the functions changes, or the parameters of any of the functions along the call chains change, the transaction will need to be fully re-executed. Styx in that case will have to re-schedule that transaction in the next epoch.
@@ 6.4. Early Commit Replies via Determinism
Implementing Styx as a fully deterministic dataflow system offers a set of advantages involving the ability to communicate transaction commits to external systems (e.g., the client), even before the state snapshots are persisted to durable storage. A traditional transactional system can respond to the client only when the requested transaction has been committed to a persistent, durable state or the write-ahead log is flushed and replicated. In Styx’s case, that would mean when an asynchronous snapshot completes (i.e., is persisted to stable storage such as S3), leading to high latency.
Since Styx implements a deterministic transactional protocol executing an agreed-upon sequence of transactions among the workers, after a failure, the system would run the same transactions with exactly the same effects. This determinism enables Styx to give early commit replies: the client can receive the reply even before a persistent snapshot is stored. The assumption here is that the message broker that persists the client requests per partition will provide to the Styx sequencers the transactions in the same order after replay as under the system’s normal operation prior to the failure, a guarantee that is typically provided by most modern message brokers including Apache Kafka that Styx currently uses.

### v4

@@ 4. Styx’s Architecture
In this section, we describe the components (Figure 3) and the main design decisions of Styx.
@@ 4.1. Components
Coordinator. The coordinator manages and monitors Styx’s workers, as well as the runtime state of the cluster (transactional metadata, dataflow state, partition locations, etc.). It also performs scheduling and health monitoring. Styx monitors the cluster’s health using a heartbeat mechanism and initiates the fault-tolerance mechanism (Section 7) once a worker fails.
@@
Worker. As depicted in Figure 3, the worker is the primary component of Styx, processing transactions, receiving or sending remote function calls, and managing state. The worker consists of two primary coroutines. The first coroutine ingests messages for its assigned partitions from a durable queue and sequences them. The second coroutine receives a set of sequenced transactions and initiates the transaction processing. By utilizing the coroutine execution model, Styx increases its efficiency since the most significant latency factor is waiting for network or state-access calls. Coroutines allow for single-threaded concurrent execution, switching between coroutines when one gets suspended during a network call, allowing others to make progress. Once the network call is completed, the suspended coroutine resumes processing.
@@
Partitioning Stateful Entities Across Workers. Styx makes use of the entities’ key to distribute those entities and their state across a number of workers. By default, each worker is assigned a set of keys using hash partitioning.
@@
Input/Output Queue. For fault tolerance, Styx assumes a persistent input queue from which it receives requests from external systems (e.g., from a REST gateway API). Styx requires the input queue to be able to deterministically replay messages based on an offset when a failure occurs. As we detail in Section 7, the replayable input queue is necessary for Styx to produce the same sequence of transactions after the recovery is complete and to enable early commit-replies (Section 6.4). In the same way, Styx sends the result of a given transaction to an output queue from which an external system (e.g., the same REST gateway API) can receive it. Currently, Styx leverages Apache Kafka (Kreps et al., 2011).
@@
Durable Snapshot Store. Alongside the replayable queue, durable storage is necessary for storing the workers’ snapshots. Currently, Styx uses Minio, an open-source S3 clone, to store the incremental snapshots as binary data files.
4.2. Transaction Execution Pipeline
@@
Styx employs an epoch-based transactional protocol that concurrently executes a batch of transactions in each epoch. A transaction may include multiple functions that, during runtime, form a call-graph of function invocations. Each function may mutate its entity’s state, and the effects of function invocations are committed to the system state in a transactional manner. In Figure 3, once make_reservation enters the system, it is persisted and replicated by the input queue. Then, a worker ingests the call into its local sequencer that assigns a Transaction ID (TID) and processes all the encapsulated function calls as a single transaction. In the make_reservation case, the transaction consists of two functions: reserve_hotel and reserve_flight. For this example, let us assume that reserve_hotel is a local function call and reserve_flight runs on a remote worker. reserve_hotel will execute locally in an asynchronous fashion using coroutines and apply state changes. In contrast, reserve_flight will execute asynchronously on a remote worker, applying changes on the remote state.
@@ 5. Sequencing & Function Execution
The deterministic execution of functions with serializable guarantees requires a sequencing step that assigns a transaction ID (TID), which, in combination with the read/write (RW) sets, can be used for conflict resolution (Section 6). The challenge we tackle in this section is determining the boundaries of transactions (i.e., when a transaction’s execution starts and finishes), which emerges from the execution of arbitrary function call-graphs Section 5.3.
5.1. Transaction Sequencing
@@
In this section, we discuss the sequencing mechanism ( ) of Styx. Deterministic databases ensure the serializable execution of transactions by forming a global sequence. In Calvin (Thomson et al., 2012), the authors propose a partitioned sequencer that retrieves the global sequence by communicating across all partitions, performing a deterministic round-robin.
@@
Eliminating Sequencer Synchronization. Instead of the original sequencer of Calvin that sends messages for the deterministic round-robin, Styx adopts a method similar to the one followed by Mencius (Mao et al., 2008), allowing Styx to acquire a global sequence without any communication between the sequencers (). This is achieved by having each sequencer assign unique transaction identifiers (TIDs) as follows:
@@
where is the sequencer id assigned by the Styx coordinator in the registration phase, is a local counter of each sequencer specifying how many TIDs it has assigned thus far and is the total number of sequencers in the Styx cluster. In the example of Figure 4, the sequencers of the three workers will sequence , and to , and respectively. Figure 5 illustrates how those TIDs are generated in parallel. Note that, conceptually, Styx implements a partitioned sequencer where the global sequence is the union of all partitioned sequences.
Mitigating Sequence Imbalance. In case a single sequencer receives more traffic than other sequencers, its local counter () will increase more than the local counter of the rest of the sequencers. As a result, in the next epoch, sequencer would produce larger TIDs than the rest of the sequencers. This means that new transactions arriving at a less busy sequencer will receive higher priority for execution: transactions with higher TID receive less priority in our transactional protocol. In case of high contention in the workload, this would increase latencies for the busy () worker node. To avoid this, at the end of an epoch, the coordinator calculates the maximum () and communicates it to all workers so that they can adjust their local counter re-balancing sequences in every epoch. Balancing the workers’ transaction priorities reduces the 99th percentile latency.
@@
Replication and Logging. There is no need to replicate and log the sequence within Styx since the input is logged and replicated within the replayable queue. In case of failure, after transaction replay the sequencers will produce the exact same sequence (section 7.2).
5.2. Call-Graph Discovery
@@
After sequencing, Styx needs to execute the sequenced transactions and determine their call-graphs and RW sets ( ). To this end, the function execution runtime ingests a given sequence of transactions to process in a given epoch. The number of transactions per epoch is either set by a time interval (by default, one millisecond) or by a configurable maximum number of transactions that can run per epoch (by default, 1000 transactions per epoch). We have chosen an epoch-based approach since processing the incoming transactions in batches increases throughput.
@@
Styx’s runtime executes all the sequenced transactions on a snapshot of the data to discover the read/write sets. Transactions that span multiple workers will implicitly change the read/write sets of the remote workers via function calls. There is an additional issue related to discovering the RW set of a transaction: before the functions execute, the call-graph of the transaction is unknown. This is an issue because the protocol requires all transactions to be completed before proceeding to the next phase. To tackle this problem, Styx proposes a function acknowledgment scheme explained in more detail in Section 5.3.
@@
After this phase, all the stateful functions that comprise transactions will have finished execution, and the RW sets will be known. In Figure 4, transactions , , and will execute and create the following RW sets: , and and .
5.3. Function Execution Acknowledgment
@@
In the SFaaS paradigm, the call-graph formed by a transaction is unknown; functions could be coded by different developer teams and can form complex call-graphs. This uncertainty complicates determining when a transaction has completed processing, which is essential because phase can only start after all transactions have finished processing. To that end, each asynchronous function call of a given transaction is assigned an ack_share. A given function knows how many shares to create by counting the number of asynchronous function calls during its runtime. The caller function then sends the respective acknowledgment shares to the downstream functions. For instance, in Figure 6, the transaction entry-point (root of the tree) calls three remote functions, splitting the ack_share into three parts (3 x ). The left-most function invokes only one other function and passes to it its complete ack_share (). The middle function does not call any functions, so it returns the share to the root function when it completes execution, and the right-most function calls two other functions, splitting its share () to 2 x . After all the function calls are complete, the root function should have collected all the shares. When the sum of the received shares adds to 1, the root/entry-point function can safely deduce that the execution of the complete transaction is complete.
@@
This design is devised for two reasons: i) if every participating function just sent an ack when it is done, the root would not know how many acks to expect in order to decide whether the entire execution has finished, and ii) if we used floats instead of fractions we could stumble upon a challenge related to adding floating point numbers. For instance, if we consider floating-point numbers in the example mentioned above of the three function calls, the sum of all shares would not equal 1, but 0.99 since each share contributes 0.33. Subsequently, we cannot accurately round inexact division numbers; therefore, Styx uses fraction mathematics instead.
@@
A solution close to the ack_share is the one of distributed futures (Wang et al., 2021). However, it would not work in the SFaaS context as it either requires information about the entire call-graph for it to work asynchronously or it would need to create a chain of futures that would make it synchronous. Hence, it would introduce high latency for our use case.
@@ 6. Committing Transactions
After completing an epoch’s call-graph discovery, Styx needs to determine which transactions will commit and which will abort based on the transactions’ Read/Write (RW) sets and TIDs. To this end, this section presents two different commit phases: an optimistic lock-free phase that commits only the non-conflicting transactions, and a lock-based phase that only commits the transactions that were not able to commit in the first phase. The lock-based commit phase commits all conflicting transactions by acquiring locks in a TID-ordered sequence. To make the second phase faster, we have devised a caching scheme that can reuse the already discovered call-graph to avoid re-executing long function chains whenever possible (section 6.3).
@@ 6.1. Lock-free Commit Phase
In case of conflict (i.e., a transaction writes a key that another transaction also reads or writes on) similarly to (Lu et al., 2020), only the transaction with the lowest transaction ID will succeed to commit ( ). The transactions that have not been committed are put in a queue to be executed in the next phase (maintaining their previously assigned ID).
In addition, workers () send their local conflicts to every other worker through the coordinator ( messages): this way, every worker retains a global view of all the aborted/rescheduled transactions and can decide, locally, which transactions can be committed. Finally, note that transactions can also abort, not because of conflicts, but due to application logic causes (e.g., by throwing an exception). In that case, Styx removes the related entries from the read/write sets to reduce possible conflicts further.
In this phase, all the transactions that have not been part of a conflict apply their writes to the state, commit, and reply to the clients. In the example shown in Figure 4, only can commit in and due to conflicts in the RW sets of regarding and ; more specifically at keys and .
@@ 6.2. Lock-based Commit Phase
In the previous phase, , only transactions without conflicts can be committed. We now explain how Styx deals with transactions that have not been committed in a given epoch due to conflicts ( ). First, Styx acquires locks in a given sequence ordered by transaction ID. Then, it reruns all transactions concurrently since all the read/write sets are known and commits them. However, if a transaction’s read/write set changes in this phase, Styx aborts the transaction and recomputes its read/write set in the next epoch. Now, in Figure 4, can sequentially acquire locks for and , leading to their commits in and .
@@ 6.3. Call-Graph Caching
As depicted in Figure 4, the lock-based commit phase is used to execute any transactions that did not commit during the lock-free commit phase . By the time the lock-based commit phase starts, the state of the database may have changed since the lock-free commit. As a result, function invocations need to be re-executed to account for the data updates.
On the left part of Figure 7, we depict such a function invocation. At time , F1 is invoked, which in turn invokes two function chains: and . Once the two function chains finish their execution (on time and respectively), they can acknowledge their termination to the root call .
Potential for Caching. During our early experiments, we noticed cases where is invoked and the parameters with which it calls (and in turn the invocations across the call chain) do not change. The same applied to the RW set of those function invocations; the RW sets remained unchanged. Since Styx tracks those call parameters as well as the functions’ RW sets, it can cache input parameters during the lock-free commit phase and reuse them during the lock-based commit, avoiding long sequential re-executions along the call chains. This case is depicted on the right part of Figure 7: the function-call chain does not need to be invoked in a sequential manner from all the way to , leading to high latency. Instead, the individual workers can re-invoke those function calls locally and concurrently. As a result, all functions can execute in parallel and save on latency and network overhead ( in Figure 7). Furthermore, caching does not require user input, is transparent to the API, and does not depend on the synchronous or asynchronous specification. Nonetheless, synchronous calls can be automatically transformed into asynchronous ones under certain conditions (Beillahi et al., 2022; Psarakis et al., 2023).
Conditions for Parallel Function Re-invocation. Intuitively, if the parameters with which, e.g., is called, and the RW set of remains the same, we can safely assume that function can be invoked concurrently without having to be invoked sequentially by . If those functions are successfully completed and acknowledge their completion to the root function , it means that the transaction can be committed. To the contrary, if the RW set of any of the functions changes, or the parameters of any of the functions along the call chains change, the transaction must be fully re-executed. In that case, Styx will have to reschedule that transaction in the next epoch.
@@ 6.4. Early Commit Replies via Determinism
Implementing Styx as a fully deterministic dataflow system offers a set of advantages involving the ability to communicate transaction commits to external systems (e.g., the client), even before the state snapshots are persisted to durable storage. A traditional transactional system can respond to the client only when the requested transaction has been committed to a persistent, durable state or the write-ahead log is flushed and replicated. In Styx’s case, that would mean when an asynchronous snapshot completes (i.e., is persisted to durable storage such as S3), leading to high latency.
Since Styx implements a deterministic transactional protocol executing an agreed-upon sequence of transactions among the workers, after a failure, the system would run the same transactions with exactly the same effects. This determinism enables Styx to give early commit replies: the client can receive the reply even before a persistent snapshot is stored. The assumption here is that the input queue, persisting the client requests, will provide to Styx’s sequencers the requests in the same order after replay, a guarantee that is typically provided by most modern message brokers. Performing state mutations and message passing before persistence has also been explored in DARQ’s speculative execution (Li et al., 2023).